OpenStudy (anonymous):

How would a person graph y=1/2(x+4)^2-5 and find the foci and focus?

OpenStudy (anonymous):

right but how does a person get -4, -0/2?

OpenStudy (anonymous):

-4, -9/2?

OpenStudy (anonymous):

on the websouce is it given as a result

OpenStudy (anonymous):

?

OpenStudy (anonymous):

okay the standard equation is given by \[(x-h)^2=4p(y-k)\] lets convert your equation into the standard form \[y=\frac{1}{2}(x+4)^2-5\Rightarrow (y+5) = \frac{1}{2}(x+4)^2\Rightarrow 2(y+5)=(x+4)^2\] Now we have the equations looking alike \[(x-h)^2=4p(y-k)\] \[(x+4)^2=2(y+5)\] \[-h=4 \rightarrow h=-4 \text{ and 4p = 2 so p = }\frac{1}{2}\text{ and -k=5 so k = -5}\] \[\text{The focus is given by (h,k+p)}\Rightarrow \text{(-4, -5+}\frac{1}{2}\text{)}\Rightarrow (-4, -\frac{9}{2})\] hope that helps

OpenStudy (anonymous):

makes sense thanx!

OpenStudy (anonymous):

p is the directrix, directrix is line that runs perpendicular to axis of symmetry of parabola distance to focus equals distance to directrix <--- definition of parabola